OpenStudy (anonymous):

look for the intercept point between y=12-x^2 and y=x^2-6

OpenStudy (anonymous):

I assume you mean intersect? 12-x^2=x^2-6 18=2x^2 9=x^2 x=3 ,-3

OpenStudy (anonymous):

yes